minutes from weekly RHIC Planning Meeting

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Sun Apr 08 2001 - 11:35:55 EDT

  • Next message: Trine Tveter: "TPM1 dN/dEta at Eta=0 - update from Oslo"

    Dear Collegue
    Enclosed are the minutes from the weekly planning meeting
    
    Flemming
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    ----- Original Message -----
    From: "Tom Ludlam" <ludlam@bnl.gov>
    To: "Marx, Michael" <marx@sbhep.physics.sunysb.edu>; "Aronson, Sam"
    <aronsons@bnl.gov>; "Ozaki, Satoshi" <ozaki@bnl.gov>; "Bond, Peter"
    <bond1@bnl.gov>; "Hallman, Tim" <tjhallman@bnl.gov>; "Zajc, Bill"
    <zajc@nevis1.nevis.columbia.edu>; "Paul, Peter" <ppaul@bnl.gov>; "Kirk, Tom"
    <tkirk@bnl.gov>; "Christie, Bill" <christie@bnl.gov>; "Peggs, Steve"
    <peggs@bnl.gov>; "Butler, Mike" <butler1@bnl.gov>; "Busza, Wit"
    <busza@mit.edu>; "Videbaek, Flemming" <videbaek@hi2>; "Harris, John"
    <Harris@star.physics.yale.edu>; "Lowenstein, Derek"
    <lowenstein@bnldag.ags.bnl.gov>; "Baker, Mark" <mdbaker@bnl.gov>; "Yousef
    Makdisi" <makdisi@bnl.gov>; "Murtagh, Michael" <murtagh@bnl.gov>; "Wlodek
    Guryn" <guryn@bnl.gov>; "Sebastian White" <white1@bnl.gov>; "Marion
    Heimerle" <heimerle@bnl.gov>; "Robert Pak" <pak@bnl.gov>; "roser1"
    <roser1@bnl.gov>; "waldo" <waldo@bnl.gov>; "dejan" <dejan@bnl.gov>;
    "brennan" <brennan@bnl.gov>; "drees" <drees@bnl.gov>; "pilat"
    <pilat@bnl.gov>; "beavis" <beavis@bnl.gov>; "pile" <pile@bnl.gov>; "throwe"
    <throwe@bnl.gov>; "Gerry Bunce" <bunce@bnl.gov>; "Mary Campbell"
    <maryc@bnl.gov>; "Russell Betts" <betts@uic.edu>; "frawley"
    <frawley@nucott.physics.fsu.edu>; "SATOGATA" <SATOGATA@bnl.gov>
    Cc: "Steadman, Steve" <Stephen.Steadman@science.doe.gov>
    Sent: Sunday, April 08, 2001 11:29 AM
    Subject: RHIC Planning Meeting
    
    
    > Dear All,
    >
    > THE NEXT RHIC MACHINE/DETECTOR PLANNING MEETING:
    >
    > WEDNESDAY, APRIL 11,  1:30
    > BUILDING 911 LG. CONF. ROOM
    >
    > Below are the minutes from last week's meeting.
    >
    > Tom
    >
    >
    ............................................................................
    .............
    >
    > RHIC Machine/Detector Planning Meeting
    > April 4, 2001
    >
    > Thomas Roser described the current status and schedule for the
    > cool-down:
    > Helium leaks have been isolated in the 4:00 and 6:00 valve boxes.
    > Repairs will delay the beginning of the cool-down by ~1 week.  The
    > schedule for closing up the ring will not be shifted (shield walls up
    > May 1).
    > The beam energy for the heavy ion run will be 100 GeV [sqrt(s) = 200].
    > This will not exactly coincide with the proton (spin) beam energy of
    > 100.5, but it is decided not to push the DX magnets beyond the design
    > maximum for Au running.
    >
    > Dejan Trbojevic announced that the plan and schedule for RHIC start-up
    > can be found at www.rhichome.bnl.gov/AP/RHIC2001/.
    >
    > Fulvia Pilat outlined the general plans for machine studies.  It is
    > expected that once we reach stable operations, there will be a block of
    > 12 hr/week scheduled for these studies.
    >
    > Angelika Drees discussed some issues at the interface between machine
    > and experiment operations.  She needs a point-of-contact from each
    > experiment to provide details of magnetic field measurements, the
    > configuration of signals sent to and from MCR, etc.
    > It was noted that we will start up with a full 60-bunch RF pattern--
    > thus, if we run with a small number of filled buckets, they will not be
    > equally spaced around the machine (as was the case last year), and so
    > there may not be collisions in every area.
    > All experiments expressed a desire for a scheduled weekly maintenance
    > shift during extended periods of stable beam running.
    >
    > PHENIX (Tony Frawley):  The South Muon Arm is in place.  The repaired
    > Drift Chamber was re-installed this week.  PHENIX should be ready to
    > start commissioning by April 19.
    >
    > STAR (Bill Christie):  The detector components for the upcoming run are
    > all in place.  Electronics for the readout of EMC modules has not yet
    > arrived at BNL-- hope to have some of the 24 modules (1/5 of full
    > barrel) instrumented as the run begins.  The shield wall will be up on
    > May 1.  The magnet pole tips will be moved into place shortly
    > thereafter.  Critical path item for the start-up: completion and
    > certification of safety interlocks.
    >
    > BRAHMS (Dana Beavis):  The forward spectrometer is being prepared for an
    > innermost stop at 2.3 degrees.  This will require moving the beam pipe 1
    > cm off axis.  The new electrical sustation is in place, allowing the
    > spectrometer magnets to reach full field.
    >
    > pp2pp (Wlodek Guryn):  Preparation is underway for an engineering run
    > during the proton commissioning period this summer.
    >
    > PHOBOS (Russell Betts):  The Second Spectrometer Arm is in place.  All
    > systems are ready to go.
    >
    > Zero-Degree Calorimeters (Sebastian White):  The ZDCs at all four
    > experiments are being prepared, with upgraded phototube bases, to
    > provide uniformly calibrated luminosity monitoring.  Analysis work is in
    > progress to determine absolute cross sections and luminosity from the
    > ZDC measurements.
    >
    



    This archive was generated by hypermail 2b29 : Sun Apr 08 2001 - 11:35:46 EDT